# A Cost-Efficient FPGA Implementation of Tiny Transformer Model using Neural ODE

### v1

@@ Abstract
Transformer is an emerging neural network model with attention mechanism. It has been adopted to various tasks and achieved a favorable accuracy compared to CNNs and RNNs. While the attention mechanism is recognized as a general-purpose component, many of the Transformer models require a significant number of parameters compared to the CNN-based ones. To mitigate the computational complexity, recently, a hybrid approach has been proposed, which uses ResNet as a backbone architecture and replaces a part of its convolution layers with an MHSA (Multi-Head Self-Attention) mechanism. In this paper, we significantly reduce the parameter size of such models by using Neural ODE (Ordinary Differential Equation) as a backbone architecture instead of ResNet. The proposed hybrid model reduces the parameter size by 94.6% compared to the CNN-based ones without degrading the accuracy. We then deploy the proposed model on a modest-sized FPGA device for edge computing. To further reduce FPGA resource utilization, we quantize the model following QAT (Quantization Aware Training) scheme instead of PTQ (Post Training Quantization) to suppress the accuracy loss. As a result, an extremely lightweight Transformer-based model can be implemented on resource-limited FPGAs. The weights of the feature extraction network are stored on-chip to minimize the memory transfer overhead, allowing faster inference. By eliminating the overhead of memory transfers, inference can be executed seamlessly, leading to accelerated inference. The proposed FPGA implementation achieves 12.8$\times$ speedup and 9.21$\times$ energy efficiency compared to ARM Cortex-A53 CPU.
@@ 1 Introduction
Transformer-based [1] model with a sophisticated attention mechanism has been an active research topic. ViT (Vision Transformer) [2] is the first pure Transformer-based architecture for visual tasks and achieves promising results compared to existing CNN-based models. However, training such models require large-scale datasets, e.g., JFT-300M [3], limiting the application on resource-limited platforms. Generally, this is due to the lack of inductive biases (e.g., locality and translational invariance) in the attention mechanism. Thus, one of the solutions for alleviating the dependence on large datasets is to combine both convolution and attention mechanism [4, 5, 6, 7, 8, 9]. Such hybrid models can achieve the state-of-the-art accuracy for small- or medium-scale datasets in a variety of tasks.
Among these models, we focus on BoTNet (Bottleneck Transformer Network) [7], which replaces spatial convolutions in the last three bottleneck blocks in ResNet with the global self-attention and outperforms baseline models by a large margin. It achieves 84.7% top-1 accuracy on ImageNet benchmark.
Based on the growing number of IoT devices, the demand for deploying DNNs to edge devices (e.g., self-driving cars, security cameras, and drones) has increased. FPGA is becoming an attractive platform to accelerate DNNs due to its energy efficiency, reconfigurability, and short development period. We aim to implement the proposed model on a modest-sized FPGA such as Xilinx ZCU104. In our previous work [10], an MHSA mechanism was implemented on the FPGA; since a large part of the model was still performed on the host CPU, only a moderate performance improvement was achieved. In contrast, this paper implements the entire feature extraction network of the proposed model on the FPGA to fully harness its computational capability. Additionally, to reduce FPGA resources, in this paper we choose to employ a lightweight quantization technique by considering the characteristics of Neural ODE. The parameter reduction and model quantization play an important role to save the scarce on-chip memory and map the Transformer-based model onto such FPGAs.
@@
The rest of this paper is organized as follows. Sec. 2 presents related works and compares our work with them. Sec. 3 provides the background on MHSA mechanism, Neural ODE, and quantization. Sec. 4 presents the design of the proposed Transformer-based lightweight model by using Neural ODE. Sec. 5 implements the proposed model on the FPGA board using quantization technique to accelerate the inference speed. Sec. 6 shows evaluation results of the proposed model in terms of parameter size and accuracy and those of the FPGA implementation in terms of resource utilization, execution time, and energy efficiency. Sec. 7 concludes this paper and discusses our future work.
@@ 2.1 CNN and Transformer
CNNs have strong inductive biases such as spatial locality and translational invariance. The spatial locality is an assumption that each pixel only correlates to its spatially neighboring pixels. While this assumption allows to efficiently capture local structures in image recognition tasks, it also comes with a potential issue of lower performance bound. On the other hand, an attention mechanism omits such inductive biases and aims to capture global structures. In this case, the model has to be trained with large datasets to capture the important relationships between different parts of each input image, which may or may not be spatially close to each other.
@@
Many attention variants have been proposed to address the high computational cost of MHSA, which grows with the square of input size $N$. In Reformer [14], LSH (Locality Sensitive Hashing) technique is employed to efficiently retain regions with high attention scores, thereby reducing the computation complexity from $O(N^{2})$ to $O(N\mathrm{log}N)$. Other methods propose the Transformer with linear complexity $O(N)$. In linear Transformer [15], the softmax operator is decomposed into two terms by employing a kernel method. Linformer [16] is a low-rank factorization-based method. Furthermore, in Flash Attention [17], memory-bound Transformer processing is accelerated by introducing an IO-aware algorithm on GPUs. These methods are proposed to overcome the computation cost associated with lengthy texts for NLP (Natural Language Processing) tasks. A similar challenge arises in the context of high-resolution image recognition tasks. Swim Transformer [18] employs fixed-pattern and global memory-based methods to reduce the computation cost from $O(N^{2})$ to $O(N)$.
@@ 2.2 Reduction of model computational complexity
Several extensions to Neural ODE have been proposed in the literature. Neural SDE (Stochastic Differential Equation) [19] incorporates various common regularization mechanisms based on the introduction of random noise. This model shows greater robustness to input noise, whether they are intentionally adversarial or not, compared to Neural ODE. In [20], training cost is significantly reduced without sacrificing accuracy by applying two appropriate regularizations to Neural ODE. In [21], by considering the regularization, TisODE (Time-invariant steady Neural ODE) is proposed to further enhance the robustness of vanilla Neural ODE. In [22], inherent constraints in Neural ODE are identified, and more expressive model called ANODE (Augmented Neural ODE) is proposed. These approaches are crucial in achieving robustness, approximation capabilities, and performance improvements in the context of Neural ODE. While this paper incorporates Transformer architecture into a model with vanilla Neural ODE as a backbone, the same idea can be applied to these Neural ODE extensions as well, which is not the focus of this paper. To further enhance our proposed model performance, we can incorporate the above techniques as future work.
@@ 2.3 FPGA implementation of Transformer architecture
There is ongoing research on FPGA implementation of Transformer to improve the inference speed. In [23], an FPGA acceleration engine that employs a column balanced block-wise pruning is introduced for customizing the balanced block-wise matrix multiplications. FTRANS [24] uses a cyclic matrix-based weight representation for Transformer model for NLP. VAQF [25] proposes an automated method to implement a quantized ViT on Xilinx ZCU102 FPGA board that meets a target accuracy. Compared to our proposal, these works do not utilize CNNs (i.e., using pure Transformer models) and thus suffer from the drawbacks of attention mechanism (e.g., huge training cost). In addition, they consider the acceleration of large-scale Transformer models (e.g., ViT) for NLP tasks, which often require high-end FPGAs due to the large number of parameters. We instead present a tiny Transformer model with less than one million parameters that fits within a modest-sized FPGA (e.g., ZCU104) for image classification tasks. The parameters are stored on-chip to eliminate the memory transfer overhead, allowing efficient inference.
In [10], a lightweight Transformer model with Neural ODE is proposed along with its FPGA-based design. Since only an MHSA block is accelerated and the other part including Neural ODE blocks is still executed on a host CPU, the design shows only a limited performance gain. Contrary to that, in this paper, we implement the entire feature extraction network to further improve the inference speed and fully benefit from the computing power of FPGAs. Additionally, in [10], the model is quantized by simply converting inputs/outputs and weights from floating- to fixed-point representation. However, it is shown that performing such simple type conversions throughout the entire model would result in a significant degradation of accuracy [10]. In this paper, we instead employ QAT, which allows more aggressive bit-width reduction (e.g., 4-bit quantization) to save the on-chip memory while mitigating the risk of significant accuracy loss.
@@ 4 Design of the proposed model
This section describes the design of the proposed model. Specifically, we outline key design aspects of the proposed model as follows: We (1) first review the ResNet architecture and (2) introduce ODENet as a memory-efficient alternative to ResNet. We then (3) apply the DSC (Depth-wise Separable Convolution) to ODENet to further reduce parameters. Analogous to BoTNet, we (4) incorporate MHSA into ODENet to derive a hybrid architecture, and (5) apply the three modifications to MHSA to make it more hardware-friendly while preserving the accuracy.
(1) ResNet consists of pre-processing, post-processing, and multiple stages for feature extraction. Pre-processing converts input images into feature maps, and post-processing transforms feature maps into a vector with elements equal to the number of classes. Each stage comprises several ResBlocks with the same structure. Specifically, each ResBlock contains two sets of convolution (Conv) layer, batch normalization (BatchNorm) layer, and activation function (e.g., ReLU), which are cascaded sequentially. In each stage, the first ResBlock is employed for downsampling, denoted as “DS”, followed by a stack of standard ResBlocks. The ResBlock for DS halves the width and height of a feature map while doubling the number of channels, i.e., it takes the feature map of size $(Ch,H,W)$ as input and produces an output of size $(2Ch,H/2,W/2)$. Note that the ResBlock for DS and standard ResBlock only differ in the configuration of the first convolution layer. For example, a straightforward approach to achieve DS is to set the stride and output channels of the first convolution in the block to 2 and $2Ch$, respectively.
@@
(3) Instead of the standard convolution, the lightweight CNN-based models [28, 29] employ DSC, which performs convolution in the spatial and channel dimensions separately. While the standard convolution uses a total of $NMK^{2}$ parameters, DSC only uses $NK^{2}+NM$ parameters, where $N,M,$ and $K$ denote the number of input channels, the number of output channels, and kernel size, respectively. Assuming $N,M\gg K$, DSC achieves approximately $K^{2}$ times parameter size reduction. As shown in Fig. 3, our model also employs DSC in the ODEBlocks.
@@
Another modification is to use ReLU instead of softmax as an activation function. As described in Sec. 3.1, relationships between query and key (i.e., logits) are obtained by taking an inner-product of $\mathbf{Q}$ with $\mathbf{K}$, and then they are passed on to the softmax function such that the attention weights $a_{i,1},\ldots,a_{i,N}$ for the $i$-th query sum to one. An immediate advantage of this modification is that ReLU is hardware-friendly as it only consumes one comparator and one multiplexer. According to [31], the accuracy of ReLU-based attention is comparable to that of the original softmax-based one. In addition, the attention weights become sparser, which assists the analysis of information flow in the model. Besides, LayerNorm [32] is added at the end of MHSABlock to stabilize gradients and facilitate the convergence.
@@ 5 Implementation
We first discuss the FPGA implementation of the proposed model described in Sec. 4. The proposed system takes a software-hardware co-design approach. The PL (Programmable Logic) part of the FPGA receives pre-processed feature maps from the PS (Processing System) part. Subsequently, the PL part processes these feature maps and returns the output to the PS part. Then, the PS part performs image classification by running the post-processing part. Specifically, the feature extraction, i.e., the forward propagation from ODEBlock1 to MHSABlock (Fig. 2, right), is carried out on the PL part. The pre- and post-processing parts are performed on the PS part with ARM Cortex-A53 CPU @1.2GHz. The operating frequency of the PL part is set to 150MHz.
As described in Sec. 4, our proposed model significantly reduces parameter size, and thus essential parameters for the feature extraction part can be stored in on-chip BRAM and URAM buffers. That is, only input and output data are transferred between the PL part and DRAM during inference. In this case, because the data transfer amount is quite small, the performance of the proposed design is compute-bound. In our FPGA implementation, the model assumes the input image resolution of 96 $\times$ 96. Note that the input image size has less impact on the model’s overall size because only the final fully-connected layer is adjusted depending on the image size and the feature extraction part is not affected.
Our FPGA implementation has two modes, which can be selected by control registers. One mode involves the transfer of weights and LUTs for quantization from DRAM to the PL part. This ensures that the necessary parameters are stored in on-chip buffers. The other mode is for inference based on the trained parameters.

### v2

@@ Abstract
Transformer has been adopted to a wide range of tasks and shown to outperform CNNs and RNNs while it suffers from high training cost and computational complexity. To address these issues, a hybrid approach has become a recent research trend, which replaces a part of ResNet with an MHSA (Multi-Head Self-Attention). In this paper, we propose a lightweight hybrid model which uses Neural ODE (Ordinary Differential Equation) as a backbone instead of ResNet for 12.1$\times$ parameter reduction. For the STL10 dataset, the proposed model achieves 80.15% top-1 accuracy which is comparable to ResNet50. Then, the proposed model is deployed on a modest-sized FPGA device for edge computing. To further reduce FPGA resource utilization, the model is quantized following QAT (Quantization Aware Training) scheme instead of PTQ (Post Training Quantization) to suppress the accuracy loss. As a result, an extremely lightweight Transformer-based model can be implemented on resource-limited FPGAs. The weights of the feature extraction network are stored on-chip to minimize the memory transfer overhead, allowing faster inference. By eliminating the overhead of memory transfers, inference can be executed seamlessly, leading to accelerated inference. The proposed FPGA implementation achieves a 34.01$\times$ speedup for the backbone and MHSA parts, and it achieves an overall 9.85$\times$ speedup when taking into account software pre- and post-processing. It also achieves an overall 7.10$\times$ higher energy efficiency compared to the ARM Cortex-A53 CPU.
@@ 1 Introduction
Transformer-based [1] model with a sophisticated attention mechanism has been an active research topic. ViT (Vision Transformer) [2] is the first pure Transformer-based architecture for visual tasks and achieves promising results compared to existing CNN-based [3] models. However, training such models require large-scale datasets, e.g., JFT-300M [4], limiting the application on resource-limited platforms. Generally, this is due to the lack of inductive biases (e.g., locality and translational invariance) in the attention mechanism. Thus, one of the solutions for alleviating the dependence on large datasets is to combine both convolution and attention mechanism [5, 6, 7, 8, 9, 10]. Such hybrid models can achieve the state-of-the-art accuracy for small- or medium-scale datasets in a variety of tasks.
Among these models, we focus on BoTNet (Bottleneck Transformer Network) [8], which replaces spatial convolutions in the last three bottleneck blocks in ResNet [11] with the global self-attention and outperforms baseline models by a large margin. It achieves 84.7% top-1 accuracy on ImageNet benchmark.
Based on the growing number of IoT devices, the demand for deploying DNNs to edge devices (e.g., self-driving cars, security cameras, and drones) has increased. FPGA is becoming an attractive platform to accelerate DNNs due to its energy efficiency, reconfigurability, and short development period. We aim to implement the proposed model on a modest-sized FPGA such as Xilinx ZCU104. In our previous work [12], MHSA mechanism was implemented on the programmable logic part of the FPGA; since a large part of the model was still performed on the host CPU, only a moderate performance improvement was achieved. In contrast, this paper implements the entire feature extraction network of the proposed model on the FPGA to fully harness its computational capability. Additionally, to reduce FPGA resources, in this paper we choose to employ a lightweight quantization technique by considering the characteristics of Neural ODE. The parameter reduction and model quantization play an important role to save the scarce on-chip memory and map the Transformer-based model onto such FPGAs.
@@
The rest of this paper is organized as follows. Sec. 2 presents related works and compares this work with them. Sec. 3 provides the background on MHSA mechanism, Neural ODE, and quantization. Sec. 4 presents the design of the proposed Transformer-based lightweight model by using Neural ODE. Sec. 5 implements the proposed model on the FPGA board using quantization technique to accelerate the inference speed. Sec. 6 shows evaluation results of the proposed model in terms of parameter size and accuracy and those of the FPGA implementation in terms of resource utilization, execution time, and energy efficiency. Sec. 7 concludes this paper and discusses the future work.
@@ 2.1 CNN and Transformer
CNNs have strong inductive biases such as spatial locality and translational invariance. The spatial locality is an assumption that each pixel only correlates to its spatially neighboring pixels. While this assumption allows to efficiently capture local structures in image recognition tasks, it also comes with a potential issue of lower performance bound. On the other hand, the computation of the attention involves performing dot products between each pixel (i.e., all-to-all) in the feature map. Therefore, the attention mechanism omits such inductive biases and aims to capture global structures [8]. In this case, the model has to be trained with large datasets to capture the important relationships between different parts of each input image, which may or may not be spatially close to each other.
@@
Recent studies have introduced hybrid models that combine CNNs and Transformers to address specific tasks in image processing, demonstrating superior accuracy. For the task of single image deraining (i.e., removing degradations caused by rain from images), hybrid model that combines CNNs and Transformers is proposed in [16]. A hybrid network combining CNNs and Transformers is proposed for lightweight image super-resolution in [17]. For the purpose of denoising low-dose CT images, a hybrid model of CNNs and Transformers is introduced in [18]. These studies demonstrate the effectiveness of hybrid models that leverage the local feature extraction capability of CNNs and the long-range dependency modeling ability of Transformers. These hybrid approaches offer a new direction by utilizing the strengths of different models to complement each other’s weaknesses, thereby enhancing performance on specific tasks.
Many attention variants have been proposed to address the high computational cost of MHSA, which grows with the square of input size $N$. In Reformer [19], LSH (Locality Sensitive Hashing) technique is employed to efficiently retain regions with high attention scores, thereby reducing the computation complexity from $O(N^{2})$ to $O(N\mathrm{log}N)$. Other methods propose the Transformer with linear complexity $O(N)$. In linear Transformer [20], the softmax operator is decomposed into two terms by employing a kernel method. Linformer [21] is a low-rank factorization-based method. Furthermore, in Flash Attention [22], memory-bound Transformer processing is accelerated by introducing an IO-aware algorithm on GPUs. These methods are proposed to overcome the computation cost associated with lengthy texts for NLP (Natural Language Processing) tasks. A similar challenge arises in the context of high-resolution image recognition tasks. Swim Transformer [23] employs fixed-pattern and global memory-based methods to reduce the computation cost from $O(N^{2})$ to $O(N)$. By employing such methods, it is possible to reduce the computational cost of the attention mechanism without sacrificing accuracy.
@@ 2.2 Reduction of model computational complexity
Several extensions to Neural ODE have been proposed in the literature. Neural SDE (Stochastic Differential Equation) [24] incorporates various common regularization mechanisms based on the introduction of random noise. This model shows greater robustness to input noise, whether they are intentionally adversarial or not, compared to Neural ODE. In [25], training cost is significantly reduced without sacrificing accuracy by applying two appropriate regularizations to Neural ODE. In [26], by considering the regularization, TisODE (Time-invariant steady Neural ODE) is proposed to further enhance the robustness of vanilla Neural ODE. In [27], inherent constraints in Neural ODE are identified, and more expressive model called ANODE (Augmented Neural ODE) is proposed. These approaches are crucial in achieving robustness, approximation capabilities, and performance improvements in the context of Neural ODE. While this paper incorporates Transformer architecture into a model with vanilla Neural ODE as a backbone, the same idea can be applied to these Neural ODE extensions as well, which is not the focus of this paper. To further enhance the proposed model performance, the above techniques are incorporated as future work.
@@ 2.3 FPGA implementation of Transformer architecture
One of advantages of the Transformer is its ability to capture a global correlation among feature maps, thanks to the attention mechanism which can enhance the accuracy. This capability not only enhances the model accuracy but also aligns well with FPGA-based acceleration since the attention mechanism relies on matrix multiplications. This operation is inherently parallelizable as well as that in CNNs. Thus, the attention mechanism can efficiently utilize the parallel processing capability of FPGAs while improving the model accuracy compared to the traditional CNN-based implementations.
There is ongoing research on FPGA implementation of Transformer to improve the inference speed. In [28], an FPGA acceleration engine that employs a column balanced block-wise pruning is introduced for customizing the balanced block-wise matrix multiplications. FTRANS [29] uses a cyclic matrix-based weight representation for Transformer model for NLP. VAQF [30] proposes an automated method to implement a quantized ViT on Xilinx ZCU102 FPGA board that meets a target accuracy. In ME-ViT [31], ViT is implemented on Xilinx Alveo U200. It adopts a single-load policy where model parameters are loaded just once and operations are centralized within a single processing element.
2.4 Research gap and objectives
In the midst of the research trend on Transformer and CNN hybrid models [5, 6, 7, 8, 9, 10, 16, 17, 18], there are few proposals for small-scale hybrid models aimed at edge computing. The objective of this paper is to demonstrate a lightweight hybrid model of CNN and attention mechanism which can be implemented on resource-limited FPGA devices.
FPGA implementations of Transformers [28, 29, 30, 31] mentioned in Sec. 2.3 do not utilize CNNs (i.e., using pure Transformer models) and thus suffer from the drawbacks of attention mechanism (e.g., huge training cost). In addition, they consider the acceleration of large-scale Transformer models (e.g., ViT), which often require high-end FPGAs or memory access overheads to load from external memory to BRAM due to the large number of parameters. The model size is a crucial factor for deployment on resource-limited edge devices. Especially, on-chip memory is a precious resource since it can offer low-latency access while its size is limited. In this paper we thus present a tiny Transformer model with one million level parameters that fits within a modest-sized FPGA (e.g., ZCU104) for image classification tasks. The parameters are stored on-chip to eliminate the memory transfer overhead, allowing efficient inference.
In [12], a lightweight Transformer model with Neural ODE is proposed along with its FPGA-based design. Since only an MHSA block is accelerated and the other part including Neural ODE blocks is still executed on a host CPU, the design shows only a limited performance gain. Contrary to that, in this paper, we implement the entire feature extraction network to further improve the inference speed and fully benefit from the computing power of FPGAs. Additionally, in [12], the model is quantized by simply converting inputs/outputs and weights from floating- to fixed-point representation. However, it is shown that performing such simple type conversions throughout the entire model would result in a significant degradation of accuracy [12]. In this paper, QAT is instead employed, which allows more aggressive bit-width reduction (e.g., 8-bit quantization) to save the on-chip memory while mitigating the risk of significant accuracy loss.
@@ 4 Design of the proposed model
This section describes the design of the proposed model. Specifically, we outline key design aspects of the proposed model as follows: (1) the ResNet architecture is first reviewed and (2) ODENet is introduced as a memory-efficient alternative to ResNet. Then, (3) the DSC (Depth-wise Separable Convolution) is applied to ODENet to further reduce parameters. Analogous to BoTNet, (4) MHSA is incorporated into ODENet to derive a hybrid architecture, and (5) the three modifications are applied to MHSA to make it more hardware-friendly while preserving the accuracy.
(1) ResNet consists of pre-processing, post-processing, and multiple stages for feature extraction as shown in Table 1. Pre-processing converts input images into feature maps, and post-processing transforms feature maps into a vector with elements equal to the number of classes. Each stage comprises several ResBlocks with the same structure. Specifically, each ResBlock contains two sets of convolution (Conv) layer, batch normalization (BatchNorm) layer, and activation function (e.g., ReLU), which are cascaded sequentially. In each stage, the first ResBlock is employed for downsampling, denoted as “DS”, followed by a stack of standard ResBlocks. The ResBlock for DS halves the width and height of a feature map while doubling the number of channels, i.e., it takes the feature map of size $(Ch,H,W)$ as input and produces an output of size $(2Ch,H/2,W/2)$. Note that the ResBlock for DS and standard ResBlock only differ in the configuration of the first convolution layer. For example, a straightforward approach to achieve DS is to set the stride and output channels of the first convolution in the block to 2 and $2Ch$, respectively.
@@
(3) Instead of the standard convolution, the lightweight CNN-based models [34, 35] employ DSC, which performs convolution in the spatial and channel dimensions separately. While the standard convolution uses a total of $NMK^{2}$ parameters, DSC only uses $NK^{2}+NM$ parameters, where $N,M,$ and $K$ denote the number of input channels, the number of output channels, and kernel size, respectively. Assuming $N,M\gg K$, DSC achieves approximately $K^{2}$ times parameter size reduction. As shown in Fig. 3, the proposed model also employs DSC in the ODEBlocks.
@@
Another modification is to use ReLU [^1] As an activation function, ReLU6 is also tested, and it shows a comparable accuracy to ReLU. In this paper, ReLU is used as well as [37]. Using ReLU6 for higher robustness is our future work. instead of softmax as an activation function. As described in Sec. 3.1, relationships between query and key (i.e., logits) are obtained by taking an inner-product of $\mathbf{Q}$ with $\mathbf{K}$, and then they are passed on to the softmax function such that the attention weights $a_{i,1},\ldots,a_{i,N}$ for the $i$-th query sum to one. An immediate advantage of this modification is that ReLU is hardware-friendly as it only consumes one comparator and one multiplexer. According to [37], the accuracy of ReLU-based attention is comparable to that of the original softmax-based one. In addition, the attention weights become sparser, which assists the analysis of information flow in the model. Besides, LayerNorm [38] is added at the end of MHSABlock to stabilize gradients and facilitate the convergence.
@@ 5 Implementation
We first discuss the FPGA implementation of the proposed model described in Sec. 4. The proposed system takes a software-hardware co-design approach. The PL (Programmable Logic) part of the FPGA receives pre-processed feature maps from the PS (Processing System) part. Subsequently, the PL part processes these feature maps and returns the output to the PS part. Then, the PS part performs image classification by running the post-processing part. Specifically, the feature extraction, i.e., the forward propagation from ODEBlock1 to MHSABlock (Fig. 2, right), is carried out on the PL part. The pre- and post-processing parts are performed on the PS part with ARM Cortex-A53 CPU @1.2GHz. The target FPGA board is the Zynq UltraScale+ MPSoC ZCU104. The operating frequency of the PL part is set to 200MHz.
As described in Sec. 4, the proposed model significantly reduces parameter size, and thus essential parameters for the feature extraction part can be stored in on-chip BRAM and URAM buffers. That is, only input and output data are transferred between the PL part and DRAM during inference. In this case, because the data transfer amount is quite small, the performance of the proposed design is compute-bound. In the FPGA implementation, the model assumes the input image resolution of 96 $\times$ 96. Note that the input image size has less impact on the model’s parameter size because the feature map size does not affect the parameter size of convolution layers.
The FPGA implementation has two modes, which can be selected by control registers. One mode involves the transfer of weights and LUTs for quantization from DRAM to the PL part. This ensures that the necessary parameters are stored in on-chip buffers. The other mode is for inference based on the trained parameters.